On one rainy day a car starts moving with a constant acceleration of 1.2 m/s2. If a toy monkey is suspended from the ceiling of the car by a string, then at what angle the string is inclined with the vertical ? (a) tan-1 (0.25) (b) tan-1 (0.63) (d) tan- 1(0.12)(c) tan-1 (√3)

Grade:11

1 Answers

Aman gohel
6 years ago
Given,A of car=1.2m/s^2g is vertically downwards=10m/s^2Therefore,a(t/c)=a(t/e)+a(e/c)=>a(t/c)=10+(-1.2)=>Therefore, tanX=(1.2/10)=>X=tan-1(0.12)[Ans]Thus the correct option is (d)